OpenStudy (anonymous):

Fan and Medal Solve. x^2 + x – 6 = 0 A. {3, –2} B. {6, –1} C. {2, –3} D. {2, 3}

OpenStudy (anonymous):

use quadratic formula

OpenStudy (anonymous):

\[\frac{ -1 \pm \sqrt{1-(-24)} }{ 2 }=\frac{ -1 \pm 5 }{ 2 }=2,-3\]

OpenStudy (anonymous):

option C